H. Cochran, Inc., is considering a new three-year expansion project that requires an initial fixed asset investment of $2,300,000. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $2,890,000 in annual sales, with costs of $1,910,000. Assume the tax rate is 21 percent and the required return on the project is 12 percent. What is the project’s NPV? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(2,300,000/3)=$766666.667

OCF=(Sales-Costs)(1-tax rate)+Tax savings on Annual depreciation

=(2,890,000-1,910,000)(1-0.21)+(0.21*766666.667)

=$935200

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=$935200[1-(1.12)^-3]/0.12

=$935200*2.40183127

=2246192.6

NPV=Present value of inflows-Present value of outflows

=2246192.6-$2,300,000

=($53807.40)(Approx)(Negative).